Mario Brothers, a game manufacturer, has a new idea for an adventure game. It can either market the game as a traditional board game or as an interactive

DVD, but not both. Consider the following cash flows of the two mutually exclusive projects. Assume the discount rate for both projects is 10 percent.

Year Board Game DVD

0 -$1,100 -$2,500

1 670 1,650

2 800 1,370

3 190 700

A/ What is the payback period for each project?

B/ What is the NPV for each project?

C/ What is the IRR for each project

D/ What is the incremental IRR?
Answer #1

A) Board Game
Uncovered amount = Initial investment - Years before Recovery * Cash Flows = 1100-670 =430
Pay back period = Years before recovery + Cost not covered in that year/ Cash flow for that year =1+430/670 = 1.64 years

Game DVD
Uncovered amount = Initial investment - Years before Recovery * Cash Flows =2500-1650 =850
Pay back period = Years before recovery + Cost not covered in that year/ Cash flow for that year =1+850/1370 = 1.62 years

b) NPV of Board Games =-1100+670/1.1+800/1.1^2+190/1.1^3 =313.00
NPV of DVD =-2500+1650/1.1+1370/1.1^2+700/1.1^3 =658.15

c)

A B Incremental Cash flow
1 -1100 -2500 -1400
2 670 1650 980
3 800 1370 570
4 190 700 510
IRR 28.17% 26.70% 25.54%
Excel Formula IRR(A1:A4) IRR(B1:B4) IRR(C1:C4)

Incremental IRR =25.54%